# Scaling limits for fractional polyharmonic Gaussian fields

### v3

@@ 1. Introduction
Figure 1.1. Surface plots of discrete fractional polyharmonic Gaussian fields on $\Omega\coloneqq(0,1)^{2}\cap\frac{1}{100}\mathbb{Z}^{2}$ with zero boundary conditions. These discrete random functions are linearly interpolated. See also [19, Fig. 1.1] for further numerical experiments.
@@ 1.1. Fractional Gaussian Fields
Similar problems have been studied before for specific values of $s$. If $s=1$, the field is the Gaussian free field and the convergence of the discrete Gaussian free field to its continuous variant is folklore (see [26, Section 4] for related results). The proof relies on the fact that covariances of the discrete Gaussian free field can be represented using simple random walk, which in the scaling limit becomes Brownian motion. The case $10\leq s<$ 1 is addressed in [19, Section 12]222 There, a definition of the discrete FGF that is slightly different from ours is used; the proof, however, should apply to all reasonable discretizations including ours. Moreover, the scaling in [19, Section 12.2] is incorrect, as we explain in Footnote 7., and the proof of the scaling limit follows a similar strategy as for the case $s=1$, just with the $2s$-stable Lévy process taking the place of Brownian motion.
These results for $s\leq 1$ all rely on some form of a random walk representation. For $s>1$ and for our choice of boundary values, there is no such random walk representation and so proofs become much more difficult. However, if one one uses another definition of the FGF in terms of spectral powers of the ordinary Laplacian (the so-called eigenfunction FGF from [19, Section 9]), one retains a random walk representations and it is comparably easy to establish a scaling limit. On the torus, the eigenfunction FGF agrees with the ordinary FGF, and for the discrete FGF on the torus results similar to ours have been shown in [6, 9] and very recently and independently in [23]. One can also study the eigenfunction FGF on domains with boundary where it is genuinely different from the ordinary FGF. In fact, in [2], this is done not in the lattice case, but in the more complicated case of a Sierpinski gasket.
@@
Thus, previous work was restricted to $s\in[0,1]\cup\{2\}$, while our results cover the entire range $s\in[0,\infty)$. Even in the case $s\in[0,1]\cup\{2\}$, our results improve upon the previous work. Namely, the convergence in [19, Section 12] is with respect to the topology of distributions and the convergence in [8] is with respect to the topology of some negative Sobolev space (for non-optimal parameters). As an easy corollary of our result with respect to the Besov-space topology, one obtains convergence with respect to the Sobolev-topology and also with respect to the Hölder topology (both with the optimal range of parameters).
@@ 1.2. Finite difference schemes for fractional operators
The scheme for $s\leq 1$ (but general $d$) has already been studied in [15] (and the special case $d=1$ already in [17, Section 4.2] and, in more detail, in [7]) and has many desirable properties. First of all, for $s\in\mathbb{N}$, we recover the standard schemes for polyharmonic Laplacians. We also have the property that $(-\Delta_{h})^{s}(-\Delta_{h})^{s^{\prime}}=(-\Delta_{h})^{s+s^{\prime}}$. Moreover, while for other schemes the accuracy often degenerates as $s\nearrow 1$, our scheme has accuracy $h^{2}$ uniformly in $s$ (as follows from Theorem 1.3). In Remark 3.2 below, we comment on how this scheme might work in practice.
@@
However, as already mentioned in the previous subsection, our interest is more in estimates under low-regularity assumptions, i.e., in a Sobolev scale. To the best of our knowledge, such estimates are new (even in the case $d=1$, $s<1$). For the case of the Laplacian or Bilaplacian, though, such results are classical (see the textbook [18, Chapter 2] or a recent refinement for the Bilaplacian in [25]), and our result is inspired by the latter. However, the proof is quite different. Namely, the proof of [25, Theorem 2.3] relied on the Bramble–Hilbert lemma and thereby used that $s\in\mathbb{N}$. In our general setting, we use a different approach, based on the Poisson summation formula and a lengthy estimate of various error terms in Fourier space.
@@ 1.3. Main results
Let $\Theta$ be a compactly supported function with integral 1. Then $I_{h}\varphi_{h}$ converges in law with respect to the topology of $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ to $\varphi$. That is, for any $f\in\mathcal{S}(\mathbb{R}^{d})$, the random variable $(I_{h}\varphi_{h},f)_{L^{2}_{h}(h\mathbb{Z}^{d})}=\sum_{x\in h\mathbb{Z}^{d}}h^{d}I_{h}\varphi_{h}(x)f(x)$ converges in law to $(\varphi,f)_{L^{2}(\mathbb{R}^{d})}=\int_{\mathbb{R}^{d}}f(x)\varphi(x)\,\mathrm{d}x$.
@@
Several remarks are in order. First of all, a convenient example of a function satisfying (1.4) for some $k\in\mathbb{N}$ is given by the centered B-spline of order $k$ (see, e.g., [24, Chapter 4] for an introduction and [18, Section 1.9.4] for a summary in the context of finite difference schemes).
@@
Let us give some explanations regarding the linear constraints on the parameters in this result. The most important constraint is $t-s\leq 2$. It arises from the fact that the proposed finite difference scheme is of second order (see [17, Section 5.2]), so that the accurary of our scheme saturates at $h^{2}$. The condition $k>\frac{d}{2}+2s-t$ is needed in order for $\Theta_{h}\ast f$ to be continuous (so that it has a well-defined restriction to the lattice). Similarly, the condition $l>\frac{d}{2}-t$ is needed in order for $\theta_{h}\ast u$ to be continuous.
@@ 2.1. The (continuous) fractional Gaussian field
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with Lipschitz boundary and let $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ be the space of tempered distributions on $\mathbb{R}^{d}$. In what follows, we collect some results on fractional Sobolev spaces on $\Omega$. If $\Omega$ has a smooth boundary, all of them are well-known (and [29, Chapter 4] is a comprehensive reference). If $\Omega$ has merely Lipschitz boundary, the situation is slightly more complicated and we rely on the reference [30]. A brief version of some of these results is also contained in [19, Section 4.1], but some of them are not made explicit there.
@@
In [19, Section 4.2], the continuous FGF is defined as a probability measure $\mathbb{P}$ on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$. More precisely, it is defined such that when $\varphi$ is distributed according to $\mathbb{P}$, then for every Schwartz function $f\in\mathcal{S}(\mathbb{R}^{d})$ we have that $(\varphi,f)$ is a centered Gaussian with variance $\|f\|^{2}_{\dot{H}^{-s}(\Omega)}$. By [19, Theorem 2.3 and Proposition 2.4], this property defines $\mathbb{P}$ as a probability measure on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ uniquely.
@@
Let us remark that the Sobolev regularity alternatively follows from the fact the random series defining $\tilde{\varphi}$ converges in $\dot{H}^{s^{\prime}}(\Omega)$ almost surely, while the Hölder regularity also follows from [19, Proposition 6.2 and Theorem 8.3]666 N.B. There is a typo in the statement of [19, Proposition 6.2]: it should read $H-k$ instead of $H-\lceil H\rceil$.. ∎
@@ 2.2. The (discrete) fractional Gaussian field
For $s=1$ and $s=2$, respectively, this agrees (up to a possible rescaling of the lattice) with the standard definitions for the Gaussian free field [26] and the membrane model [8, 25] in the literature. For $10<s<$ 1, as already mentioned in the introduction, our definition is not the same as the one in the survey [19, Section 12] 777 We note that there are a few small errors in [19, Section 12.2]. In particular, with the definition of a discrete FGF given there, [19, Proposition 12.2] is false. The correct definition should have the density (2.2) $\exp\left(-\frac{1}{2}\,\sum_{\begin{subarray}{c}x,\,y\in\delta\mathbb{Z}^{d}\\ x\neq y\end{subarray}}C_{d,s}\frac{|f(x)-f(y)|^{2}}{|x-y|^{d+2s}}\delta^{2d}\right)$ while, in [19, Section 12.2], $\delta^{d}$ is used in place of $\delta^{2d}$. Only the definition (2.2) has the property that if we send $\delta\to 0$, we formally get the density $\exp\left(-\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}C_{d,s}\frac{|f(x)-f(y)|^{2}}{|x-y|^{d+2s}}\,\mathrm{d}x\,\mathrm{d}y\right)$ of the continuous FGF. The error in the proof of [19, Proposition 12.2] is in (12.10). The process $(Y^{\delta}_{t})_{t=0}^{\infty}$ converges to $(Y_{t})_{t=0}^{\infty}$ pathwise, and so the occupation measure of $A\cap\delta\mathbb{Z}^{d}$ of the former process converges to the occupation measure of $A$ of the latter. Thus, the additional factor of $\delta^{d}$ on the left-hand side of (12.10) is erroneous. With this error corrected, the proof of [19, Proposition 12.2] does show that the discrete FGF with density (2.2) converges in the sense of distributions. . Their definition would correspond to the operator $\widetilde{(-\Delta_{h})^{s}}$, where $\widetilde{(-\Delta_{h})^{s}}u_{h}(x)=-C_{d,s}\sum_{y\in h\mathbb{Z}^{d}\setminus\{0\}}h^{d}\frac{u_{h}(y)-u_{h}(x)}{|x-y|^{d+2s}}$. From the point of view of numerical analysis, this discretization is quite degenerate (cf. the discussion at the beginning of Section 4.4 in [17]).
@@
Nonetheless our methods apply to other discretizations of the fractional FGF as well, including the one in [19]. The basis for all our arguments is (1.3). If, instead of $O(h^{2}|\xi|^{2})$, one only had $O(h^{\kappa}|\xi|^{\kappa})$ for some $0<\kappa\leq 2$, then Theorem 1.3 would hold with the restriction $t-s\leq\kappa$. For the application towards Theorems 1.1 and 1.2, such an estimate for some $\kappa>0$ is good enough. So our results on the scaling limit of the discrete FGF also hold for other models as long as we can show a version of (1.3) with an error term $O(h^{\kappa}|\xi|^{\kappa})$. That is, we need good control over the asymptotics of the symbol of the corresponding operator. In general, it can be non-trivial to obtain these asymptotics. However, for the case of $\widetilde{(-\Delta_{h})^{s}}$, or more generally for schemes of the form $-C_{d,s}\sum_{y\in h\mathbb{Z}^{d}\setminus\{0\}}h^{d}k(|x-y|)(u_{h}(y)-u_{h}(x))$ for $k(x)=\frac{1}{|x|^{d+2s}}\left(1+O\left(\frac{1}{|x|^{\kappa^{\prime}}}\right)\right)$ for $10<s<$ 1, the relevant calculations can be done as in [16, Appendix]. In particular, also for the version of the discrete FGF from [19] we have not just Theorem 1.1 (as shown in [19, Proposition 12.2]), but also Theorem 1.2. This essentially answers the question in [12, Remark 6].
@@ 3. Rigorous estimates for the finite difference scheme
By the Poisson summation formula (see, e.g., [22, Chapter 4.4]), for any Schwartz function $f$, we have
@@
If one is willing to accept an additional error term, then a more efficient way to compute an approximation to the entries of $((-\Delta_{h})^{s})_{x,\,y\in\Omega_{h}}$ was suggested in [15]: choose a parameter $h^{\prime}\leq h$, and approximate the integral over $\left(-\frac{\pi}{h},\frac{\pi}{h}\right)^{d}$ appearing in the definition of $(-\Delta_{h})^{s}$ by a Riemann sum on a lattice of width $\frac{h^{\prime}}{h}$. The advantage is that this Riemann sum can be computed very efficiently using the fast Fourier transform. Moreover, in [15, Section 4.2], it is suggested that this should lead to an additional error of order $O\left(\frac{h^{\prime d+2s}}{h^{2s}}\right)$. In other words, if we choose $h^{\prime}\leq h^{(2s+2)/(2s+d)}$, the error should be of order $h^{2}$ and thus not bigger than the error in Theorem 1.3. While the error estimate in [15, Section 4.2] is not rigorous, it should be possible to give a full proof.
@@ 4.1. Scaling limit in the space of distributions
Moreover, let $\tilde{\Theta},\,\tilde{\theta}$ be functions satisfying the assumptions of Theorem 1.3 with $k\coloneqq\max\left(\frac{d}{2}+s,s\right)$ and $l\coloneqq\max\left(\frac{d}{2}-s,0\right)$; for example, let us take $\tilde{\Theta}$ to be a B-spline of order $\lceil k\rceil$ as in [18, Section 1.9.4] and $\tilde{\theta}$ any smooth mollifier. Then, let us define $\tilde{\Theta}_{h}$ and $\tilde{\theta}_{h}$ as before and let $\tilde{u}_{h}\colon\,h\mathbb{Z}^{d}\to\mathbb{R}$ be the solution of

### v4

@@ 1. Introduction
Figure 1.1. Surface plots of discrete fractional polyharmonic Gaussian fields on $\Omega\coloneqq(0,1)^{2}\cap\frac{1}{100}\mathbb{Z}^{2}$ with zero boundary conditions. These discrete random functions are linearly interpolated. See also [20, Fig. 1.1] for further numerical experiments.
@@ 1.1. Fractional Gaussian Fields
Similar problems have been studied before for specific values of $s$. If $s=1$, the field is the Gaussian free field and the convergence of the discrete Gaussian free field to its continuous variant is folklore (see [26, Section 4] for related results). The proof relies on the fact that covariances of the discrete Gaussian free field can be represented using simple random walk, which in the scaling limit becomes Brownian motion. The case $10\leq s<$ 1 is addressed in [20, Section 12]222 There, a definition of the discrete FGF that is slightly different from ours is used; the proof, however, should apply to all reasonable discretizations including ours. Moreover, the scaling in [20, Section 12.2] is incorrect, as we explain in Footnote 7., and the proof of the scaling limit follows a similar strategy as for the case $s=1$, just with the $2s$-stable Lévy process taking the place of Brownian motion.
These results for $s\leq 1$ all rely on some form of a random walk representation. For $s>1$ and for our choice of boundary values, there is no such random walk representation and so proofs become much more difficult. However, if one one uses another definition of the FGF in terms of spectral powers of the ordinary Laplacian (the so-called eigenfunction FGF from [20, Section 9]), one retains a random walk representations and it is comparably easy to establish a scaling limit. On the torus, the eigenfunction FGF agrees with the ordinary FGF, and for the discrete FGF on the torus results similar to ours have been shown in [6, 9] and very recently and independently in [11]. One can also study the eigenfunction FGF on domains with boundary where it is genuinely different from the ordinary FGF. In fact, in [2], this is done not in the lattice case, but in the more complicated case of a Sierpinski gasket.
@@
Thus, previous work was restricted to $s\in[0,1]\cup\{2\}$, while our results cover the entire range $s\in[0,\infty)$. Even in the case $s\in[0,1]\cup\{2\}$, our results improve upon the previous work. Namely, the convergence in [20, Section 12] is with respect to the topology of distributions and the convergence in [8] is with respect to the topology of some negative Sobolev space (for non-optimal parameters). As an easy corollary of our result with respect to the Besov-space topology, one obtains convergence with respect to the Sobolev-topology and also with respect to the Hölder topology (both with the optimal range of parameters).
@@ 1.2. Finite difference schemes for fractional operators
The scheme for $s\leq 1$ (but general $d$) has already been studied in [16] (and the special case $d=1$ already in [18, Section 4.2] and, in more detail, in [7]) and has many desirable properties. First of all, for $s\in\mathbb{N}$, we recover the standard schemes for polyharmonic Laplacians. We also have the property that $(-\Delta_{h})^{s}(-\Delta_{h})^{s^{\prime}}=(-\Delta_{h})^{s+s^{\prime}}$. Moreover, while for other schemes the accuracy often degenerates as $s\nearrow 1$, our scheme has accuracy $h^{2}$ uniformly in $s$ (as follows from Theorem 1.3). In Remark 3.2 below, we comment on how this scheme might work in practice.
@@
However, as already mentioned in the previous subsection, our interest is more in estimates under low-regularity assumptions, i.e., in a Sobolev scale. To the best of our knowledge, such estimates are new (even in the case $d=1$, $s<1$). For the case of the Laplacian or Bilaplacian, though, such results are classical (see the textbook [19, Chapter 2] or a recent refinement for the Bilaplacian in [25]), and our result is inspired by the latter. However, the proof is quite different. Namely, the proof of [25, Theorem 2.3] relied on the Bramble–Hilbert lemma and thereby used that $s\in\mathbb{N}$. In our general setting, we use a different approach, based on the Poisson summation formula and a lengthy estimate of various error terms in Fourier space.
@@ 1.3. Main results
Let $\Theta$ be a compactly supported function with integral 1. Then $I_{h}\varphi_{h}$ converges in law with respect to the topology of $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ to $\varphi$. That is, for any $f\in\mathcal{S}(\mathbb{R}^{d})$, the random variable $(I_{h}\varphi_{h},f)_{L^{2}(\mathbb{R}^{d})}=\int_{\mathbb{R}^{d}}I_{h}\varphi_{h}(x)f(x)\,\mathrm{d}x$ converges in law to $(\varphi,f)_{L^{2}(\mathbb{R}^{d})}=\int_{\mathbb{R}^{d}}\varphi(x)f(x)\,\mathrm{d}x$.
@@
Several remarks are in order. First of all, a convenient example of a function satisfying (1.4) for some $k\in\mathbb{N}$ is given by the centered B-spline of order $k$ (see, e.g., [24, Chapter 4] for an introduction and [19, Section 1.9.4] for a summary in the context of finite difference schemes).
@@
Let us give some explanations regarding the linear constraints on the parameters in this result. The most important constraint is $t-s\leq 2$. It arises from the fact that the proposed finite difference scheme is of second order (see [18, Section 5.2]), so that the accurary of our scheme saturates at $h^{2}$. The condition $k>\frac{d}{2}+2s-t$ is needed in order for $\Theta_{h}\ast f$ to be continuous (so that it has a well-defined restriction to the lattice). Similarly, the condition $l>\frac{d}{2}-t$ is needed in order for $\theta_{h}\ast u$ to be continuous.
@@ 2.1. The (continuous) fractional Gaussian field
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with Lipschitz boundary and let $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ be the space of tempered distributions on $\mathbb{R}^{d}$. In what follows, we collect some results on fractional Sobolev spaces on $\Omega$. If $\Omega$ has a smooth boundary, all of them are well-known (and [29, Chapter 4] is a comprehensive reference). If $\Omega$ has merely Lipschitz boundary, the situation is slightly more complicated and we rely on the reference [30]. A brief version of some of these results is also contained in [20, Section 4.1], but some of them are not made explicit there.
@@
In [20, Section 4.2], the continuous FGF is defined as a probability measure $\mathbb{P}$ on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$. More precisely, it is defined such that when $\varphi$ is distributed according to $\mathbb{P}$, then for every Schwartz function $f\in\mathcal{S}(\mathbb{R}^{d})$ we have that $(\varphi,f)$ is a centered Gaussian with variance $\|f\|^{2}_{\dot{H}^{-s}(\Omega)}$. By [20, Theorem 2.3 and Proposition 2.4], this property defines $\mathbb{P}$ as a probability measure on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ uniquely.
@@
Let us remark that the Sobolev regularity alternatively follows from the fact the random series defining $\tilde{\varphi}$ converges in $\dot{H}^{s^{\prime}}(\Omega)$ almost surely, while the Hölder regularity also follows from [20, Proposition 6.2 and Theorem 8.3]666 N.B. There is a typo in the statement of [20, Proposition 6.2]: it should read $H-k$ instead of $H-\lceil H\rceil$.. ∎
@@ 2.2. The (discrete) fractional Gaussian field
For $s=1$ and $s=2$, respectively, this agrees (up to a possible rescaling of the lattice) with the standard definitions for the Gaussian free field [26] and the membrane model [8, 25] in the literature. For $10<s<$ 1, as already mentioned in the introduction, our definition is not the same as the one in the survey [20, Section 12] 777 We note that there are a few small errors in [20, Section 12.2]. In particular, with the definition of a discrete FGF given there, [20, Proposition 12.2] is false. The correct definition should have the density (2.2) $\exp\left(-\frac{1}{2}\,\sum_{\begin{subarray}{c}x,\,y\in\delta\mathbb{Z}^{d}\\ x\neq y\end{subarray}}C_{d,s}\frac{|f(x)-f(y)|^{2}}{|x-y|^{d+2s}}\delta^{2d}\right)$ while, in [20, Section 12.2], $\delta^{d}$ is used in place of $\delta^{2d}$. Only the definition (2.2) has the property that if we send $\delta\to 0$, we formally get the density $\exp\left(-\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}C_{d,s}\frac{|f(x)-f(y)|^{2}}{|x-y|^{d+2s}}\,\mathrm{d}x\,\mathrm{d}y\right)$ of the continuous FGF. The error in the proof of [20, Proposition 12.2] is in (12.10). The process $(Y^{\delta}_{t})_{t=0}^{\infty}$ converges to $(Y_{t})_{t=0}^{\infty}$ pathwise, and so the occupation measure of $A\cap\delta\mathbb{Z}^{d}$ of the former process converges to the occupation measure of $A$ of the latter. Thus, the additional factor of $\delta^{d}$ on the left-hand side of (12.10) is erroneous. With this error corrected, the proof of [20, Proposition 12.2] does show that the discrete FGF with density (2.2) converges in the sense of distributions. . Their definition would correspond to the operator $\widetilde{(-\Delta_{h})^{s}}$, where $\widetilde{(-\Delta_{h})^{s}}u_{h}(x)=-C_{d,s}\sum_{y\in h\mathbb{Z}^{d}\setminus\{0\}}h^{d}\frac{u_{h}(y)-u_{h}(x)}{|x-y|^{d+2s}}$. From the point of view of numerical analysis, this discretization is quite degenerate (cf. the discussion at the beginning of Section 4.4 in [18]).
@@
Nonetheless our methods apply to other discretizations of the fractional FGF as well, including the one in [20]. The basis for all our arguments is (1.3). If, instead of $O(h^{2}|\xi|^{2})$, one only had $O(h^{\kappa}|\xi|^{\kappa})$ for some $0<\kappa\leq 2$, then Theorem 1.3 would hold with the restriction $t-s\leq\kappa$. For the application towards Theorems 1.1 and 1.2, such an estimate for some $\kappa>0$ is good enough. So our results on the scaling limit of the discrete FGF also hold for other models as long as we can show a version of (1.3) with an error term $O(h^{\kappa}|\xi|^{\kappa})$. That is, we need good control over the asymptotics of the symbol of the corresponding operator. In general, it can be non-trivial to obtain these asymptotics. However, for the case of $\widetilde{(-\Delta_{h})^{s}}$, or more generally for schemes of the form $-C_{d,s}\sum_{y\in h\mathbb{Z}^{d}\setminus\{0\}}h^{d}k(|x-y|)(u_{h}(y)-u_{h}(x))$ for $k(x)=\frac{1}{|x|^{d+2s}}\left(1+O\left(\frac{1}{|x|^{\kappa^{\prime}}}\right)\right)$ for $10<s<$ 1, the relevant calculations can be done as in [17, Appendix]. In particular, also for the version of the discrete FGF from [20] we have not just Theorem 1.1 (as shown in [20, Proposition 12.2]), but also Theorem 1.2. This essentially answers the question in [13, Remark 6].
@@ 3. Rigorous estimates for the finite difference scheme
By the Poisson summation formula (see, e.g., [23, Chapter 4.4]), for any Schwartz function $f$, we have
@@
If one is willing to accept an additional error term, then a more efficient way to compute an approximation to the entries of $((-\Delta_{h})^{s})_{x,\,y\in\Omega_{h}}$ was suggested in [16]: choose a parameter $h^{\prime}\leq h$, and approximate the integral over $\left(-\frac{\pi}{h},\frac{\pi}{h}\right)^{d}$ appearing in the definition of $(-\Delta_{h})^{s}$ by a Riemann sum on a lattice of width $\frac{h^{\prime}}{h}$. The advantage is that this Riemann sum can be computed very efficiently using the fast Fourier transform. Moreover, in [16, Section 4.2], it is suggested that this should lead to an additional error of order $O\left(\frac{h^{\prime d+2s}}{h^{2s}}\right)$. In other words, if we choose $h^{\prime}\leq h^{(2s+2)/(2s+d)}$, the error should be of order $h^{2}$ and thus not bigger than the error in Theorem 1.3. While the error estimate in [16, Section 4.2] is not rigorous, it should be possible to give a full proof.
@@ 4.1. Scaling limit in the space of distributions
Moreover, let $\tilde{\Theta},\,\tilde{\theta}$ be functions satisfying the assumptions of Theorem 1.3 with $k\coloneqq\max\left(\frac{d}{2}+s,s\right)$ and $l\coloneqq\max\left(\frac{d}{2}-s,0\right)$; for example, let us take $\tilde{\Theta}$ to be a B-spline of order $\lceil k\rceil$ as in [19, Section 1.9.4] and $\tilde{\theta}$ any smooth mollifier. Then, let us define $\tilde{\Theta}_{h}$ and $\tilde{\theta}_{h}$ as before and let $\tilde{u}_{h}\colon\,h\mathbb{Z}^{d}\to\mathbb{R}$ be the solution of